OpenStudy (anonymous):

Evaluate the exponential function for the given value of x. f(x)=(1/4)^x x=4 Simplify your answer- f(x)=____

jimthompson5910 (jim_thompson5910):

f(x)=(1/4)^x f(4)=(1/4)^4 f(4) = 1/256

OpenStudy (anonymous):

Given f(x)=(1/4)^x and x=4. or,f(4)=(1/4)^4 or,f(4)=(1/4)*(1/4)*(1/4)*(1/4) or,f(4)=1/256